# Design of supercurrent diode by vortex phase texture

### v3

@@ III.1 Single vortex configuration
Moving to even high-winding, $V^{\mathrm{L}}_{0}=2$, we find that the maximal amplitude of the rectification occurs for vortex core positions that are now closer to the lateral edges of the superconductor but still midway from the interface [Figs. 2 (f,g)]. Inspection of the harmonics content of the current phase relation confirms that the sign change of the rectification amplitude $\eta$ occurs when $J_{1}$ reverses its sign [Figs. 2 (h,i)]. Furthermore, we find that the sign change of $\eta$ does not directly follow the $I_{1}$ sign change and the maximum of the rectification ($\eta\sim 25\%$) arises for vortex core positions whose $I_{1}$, $I_{2}$ and $J_{1}$ have comparable size. Let us then consider a vortex with odd high-winding number $V^{\mathrm{L}}_{0}=3$ assuming a square ($\alpha=1$) shape for the superconducting lead [Figs. 2 (j)-(m)]. For $V^{\mathrm{L}}_{0}=3$, we observe that the maximal values of the rectification amplitude occur for the vortex core position that is away from the mirror lines of the superconducting lead [Figs. 2 (j,k)]. The rectification can reach $30\%$ amplitude and there are multiple positions of the vortex core for which $\eta$ is vanishing [Figs. 2 (j,k)].
The analysis indicates that a key element to maximize the rectification is represented by the search of vortex configurations for which the first harmonics (even and odd-parity) are concomitantly almost vanishing. This is a general rule to get large rectification [51] (see also Appendix C). To this aim, one can make an analytical analysis of the first harmonics by considering the direct process of Cooper pairs transfer across the junction as given by [52, 50]
@@
×$\times[\sin\varphi\cos\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})+\cos\varphi\sin\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})].$
This relation is general and can be applied to any type of vortex phase texture. Here, from this expression, we can deduce how nonreciprocal supercurrent directly links to the winding of the phase vortex. Focusing on the coefficients of $\sin\varphi$ and $\cos\varphi$, the structure of the vortex indeed directly impacts on the even and odd-parity components of the first harmonics of the supercurrent. For convenience and in order to primarily extract the angular dependence of the harmonics content we neglect the amplitude variation in the vortex core. Hence, the effective first harmonics components $\bar{I}_{1}(\bm{r}^{\mathrm{L}}_{0})$ and $\bar{J}_{1}(\bm{r}^{\mathrm{L}}_{0})$ of $\tilde{I}$ are substantially given by
@@
$\bar{J}_{1}(\bm{r}^{\mathrm{L}}_{0})$ $=\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\sin\varphi_{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})$ (19)
@@
We find that their nodal lines can cross in different sites depending on the winding number and aspect ratio of the superconducting lead [Fig. 3]. In particular, for the case of even winding, $V^{\mathrm{L}}_{0}=2$, the breaking of $C_{4}$ rotational symmetry for a rectangular-shaped superconductor induces a shift of the nodal line of the odd-harmonic $\bar{I}_{1}$ towards the edge of the superconductor [Fig. 3 (c)]. Then, the crossing with the nodal lines for $\bar{J}_{1}$ shifts from $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})=(-22a,0)$ to $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})\sim(-22a,\pm 10.5a)$. When considering a vortex with winding $V^{\mathrm{L}}_{0}=3$, due to the higher angular components, the crossings of the vanishing lines for $\bar{I}_{1}$ are in multiple points within the superconducting domain at $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})\sim(-7a,\pm 7.5a)$ and $(-22a,\pm 7.5a)$ [Fig. 3 (e,f)].
III.2 Two-vortex configuration
Figure 4: (a) Schematic illustration for two vortices in the left-side superconductor. (b,c) Rectification $\eta$ and (d,e) $I_{1}$, $I_{2}$, $J_{1}$ as a function of the literal direction $y^{\mathrm{L}}_{1}$ for (b,d) $(V^{\mathrm{L}}_{0},V^{\mathrm{L}}_{1})=(+1,+1)$ and (c,e) $(V^{\mathrm{L}}_{0},V^{\mathrm{L}}_{1})=(+1,-1)$. We set the core positions as $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})=(-30,0.5)$ and $x^{\mathrm{L}}_{1}=-15$ ($N^{\mathrm{L}}_{x}=N^{\mathrm{R}}_{x}=45$, $N^{\mathrm{N}}_{x}=10$, $N_{y}=30$, $z^{\mathrm{L}}_{0}=z^{\mathrm{L}}_{1}=10a$). $I_{0}=0.122|\Delta_{0}|$ stands for the maximum Josephson current without any phase vortices in superconductors.
@@ III.2 Two-vortex configuration
$\Delta_{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0},\bm{r}^{\mathrm{L}}_{1})=|\Delta_{0}|\prod^{1}_{l=0}\tilde{\Theta}_{l}(\bm{j},\bm{r}_{l})e^{i\varphi^{\mathrm{L}}_{\mathrm{v}l}(\bm{j},\bm{r}_{l})},$ (20)
@@
$\tilde{\Theta}_{l}(\bm{j},\bm{r}_{l})=\tanh\left[\frac{|\bm{j}-\bm{r}^{\mathrm{L}}_{l}|}{z^{\mathrm{L}}_{l}}\right],$ (21)
the size of vortices $z^{\mathrm{L}}_{l}$, and the core positions $\bm{r}^{\mathrm{L}}_{l}=(x^{\mathrm{L}}_{l},y^{\mathrm{L}}_{l})$ for $l=0,1$. Each phase vortex is given by
$\varphi^{\mathrm{L}}_{\mathrm{v}l}(\bm{j})=V^{\mathrm{L}}_{l}\arg[(j_{x}a-x^{\mathrm{L}}_{l})+i(j_{y}a-y^{\mathrm{L}}_{l})],$ (22)
where $V^{\mathrm{l}}$ is the number of winding for each phase vortex.
@@
Figure 5: (a,b) $\bar{I}_{1}$ and (c,d) $\bar{J}_{1}$ for the $\bm{j}^{\mathrm{L}}_{1}$ space at (a,c) $(V^{\mathrm{L}}_{0},V^{\mathrm{L}}_{1})=(+1,+1)$ and (b,d) $(V^{\mathrm{L}}_{0},V^{\mathrm{L}}_{1})=(+1,-1)$. We set the core positions as $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})=(-30a,0.5a)$ and $z^{\mathrm{L}}_{1}=10a$ ($N^{\mathrm{L}}_{x}=N^{\mathrm{R}}_{x}=45$ and $N_{y}=30$). The Gray line indicates the scanning site in Fig. 4.

### v5

@@ Introduction
Moving to even high-winding, $V^{\mathrm{L}}_{0}=2$, we find that the maximal amplitude of the rectification occurs for vortex core positions that are now closer to the lateral edges of the superconductor but still midway from the interface [Figs. 2 (f,g)]. Inspection of the harmonics content of the current phase relation confirms that the sign change of the rectification amplitude $\eta$ occurs when $J_{1}$ reverses its sign [Figs. 2 (h,i)]. Furthermore, we find that the sign change of $\eta$ does not directly follow the $I_{1}$ sign change and the maximum of the rectification ($\eta\sim 25\%$) arises for vortex core positions whose $I_{1}$, $I_{2}$, and $J_{1}$ have comparable size. Let us then consider a vortex with odd high-winding number $V^{\mathrm{L}}_{0}=3$ assuming a square ($\alpha=1$) shape for the superconducting lead [Figs. 2 (j)-(m)]. For $V^{\mathrm{L}}_{0}=3$, we observe that the maximal values of the rectification amplitude occur for the vortex core position that is away from the mirror lines of the superconducting lead [Figs. 2 (j,k)]. The rectification can reach $30\%$ amplitude and there are multiple positions of the vortex core for which $\eta$ is vanishing [Figs. 2 (j,k)]. Now, when considering the specific case of the current phase relation in the presence of a vortex, since the second harmonic components are usually smaller than the first harmonic, then the condition to become comparable in amplitude is usually met when the first harmonics are vanishingly small or change sign. This happens, e.g. nearby $x_{0}^{\mathrm{L}}=22a$ or $7a$, and in general, nearby the points where the first harmonics become vanishing. In this context, the appearance of a dip and the peak around specific vortex core positions (see Fig. 2 b) results from the simultaneous fact that the first odd-parity harmonic reverses sign and that the amplitudes of the first and second harmonics are of similar magnitude.
The analysis indicates that a key element to maximize the rectification is represented by the search for vortex configurations for which the first harmonics (even and odd-parity) are concomitantly almost vanishing. This is a general rule to get large rectification [61] (see Supplementary Info - Section D). To this aim, one can make an analytical analysis of the first harmonics by considering the direct process of Cooper pairs transfer across the junction as given by [62, 50]
@@
$\times[\sin\varphi\cos\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})+\cos\varphi\sin\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})].$
This relation is general and can be applied to any type of vortex phase texture. Here, from this expression, we can deduce how the nonreciprocal supercurrent directly links to the winding of the phase vortex. Focusing on the coefficients of $\sin\varphi$ and $\cos\varphi$, the structure of the vortex indeed directly impacts the even and odd-parity components of the first harmonics of the supercurrent. For convenience and in order to primarily extract the angular dependence of the harmonic content, we neglect the amplitude variation in the vortex core. Hence, the effective first harmonics components $\bar{I}_{1}(\bm{r}^{\mathrm{L}}_{0})$ and $\bar{J}_{1}(\bm{r}^{\mathrm{L}}_{0})$ of $\tilde{I}$ are substantially given by
@@
$\bar{J}_{1}(\bm{r}^{\mathrm{L}}_{0})$ $=\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\sin\varphi_{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0}).$ (19)
@@
We find that their nodal lines can cross in different sites depending on the winding number and aspect ratio of the superconducting lead [Fig. 3]. In particular, for the case of even winding, $V^{\mathrm{L}}_{0}=2$, the breaking of $C_{4}$ rotational symmetry for a rectangular-shaped superconductor induces a shift of the nodal line of the odd-harmonic $\bar{I}_{1}$ towards the edge of the superconductor [Fig. 3 (d),(e)]. Then, the crossing with the nodal lines for $\bar{J}_{1}$ shifts from $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})=(-22a,0)$ to $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})\sim(-22a,\pm 10.5a)$. When considering a vortex with winding $V^{\mathrm{L}}_{0}=3$, due to the higher angular components, the crossings of the vanishing lines for $\bar{I}_{1}$ are in multiple points within the superconducting domain at $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})\sim(-7a,\pm 7.5a)$ and $(-22a,\pm 7.5a)$ [Fig. 3 (g,h)]. To verify how the spatial profile of the rectification amplitude varies with different vortex states characterized by their winding numbers, we performed numerical computations of the supercurrent in the superconducting junction. These calculations reveal that the rectification pattern exhibits distinct spatial features depending on the vortex winding number. As illustrated in Fig. 3 (c), (f), and (i), the pattern of rectification amplitude changes qualitatively with different vortex states. Specifically, for the case where the vortex has a winding number $V^{\mathrm{L}}_{0}=1$, there is a single region where the rectification amplitude vanishes. This region corresponds to the mirror symmetry line with respect to the transformation $y\to-y$. The sign change of the rectification amplitude primarily occurs when crossing this horizontal mirror line at $y=0$. In contrast, for a vortex with winding number $V^{\mathrm{L}}_{0}=2$, the spatial pattern features two lines where the rectification amplitude changes sign. These sign-change lines are associated with the vortex core’s position relative to the interface, which can move from regions far from the interface to those closer to it. As the vortex winding number increases further to $V^{\mathrm{L}}_{0}=3$, the number of nodal lines—i.e., lines where the rectification amplitude vanishes—expands to four, indicating a more complex spatial structure of the supercurrent distribution influenced by the vortex’s winding number. These findings confirm the qualitative expectation based on the analysis of the first harmonics of the current phase relation.
Two-vortex configuration.
Figure 4: Space dependence of rectifications with two vortices. (a) Schematic illustration of two vortices in the left-side superconductor. SC-L, N, and SC-R mean the Left-side superconductor (SC), the Normal metal, and the Right-side SC. $\bm{r}^{\mathrm{L}}_{m}=(x^{\mathrm{L}}_{m},y^{\mathrm{L}}_{l})$ with $m=1,2$ indicate core position for each vortex. $V^{\mathrm{L}}_{0,1}$ indicate the winding number of each phase vortex. $N^{\mathrm{L}}_{x}$, $N^{\mathrm{R}}_{x}$, $N^{\mathrm{N}}_{x}$, $N_{y}$, and $a$ denote the number of sites in the left and right-side SCs, and in the normal metal along the $x$-direction, and number of sites along the $y$-direction, and lattice constant. (b,c) Rectification $\eta$ and (d,e) $I_{1}$, $I_{2}$, $J_{1}$ as a function of the literal direction $y^{\mathrm{L}}_{1}$ for each (b,d) $(V^{\mathrm{L}}_{0},V^{\mathrm{L}}_{1})=(+1,+1)$ and (c,e) $(V^{\mathrm{L}}_{0},V^{\mathrm{L}}_{1})=(+1,-1)$. We set the core positions as $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})=(-30,0.5)$ and $x^{\mathrm{L}}_{1}=-15$. $I_{0}=0.122|\Delta_{0}|(2e/\hbar)$ stands for the maximum Josephson current without any phase vortices in superconductors. We select the parameters: $|\Delta_{0}|=0.02t$ (superconducting energy gap amplitude), $t_{\mathrm{int}}=0.90$ (transparency at the interface), $N^{\mathrm{L}}_{x}=N^{\mathrm{R}}_{x}=45$, $N^{\mathrm{N}}_{x}=10$, $N_{y}=30$, and $z^{\mathrm{L}}_{0}=z^{\mathrm{L}}_{1}=10a$ (each vortex size).
@@
$\Delta_{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0},\bm{r}^{\mathrm{L}}_{1})=|\Delta_{0}|\prod^{1}_{m=0}\tilde{\Theta}_{m}(\bm{j},\bm{r}_{m})e^{i\varphi^{\mathrm{L}}_{\mathrm{v}m}(\bm{j},\bm{r}_{m})},$ (20)
@@
$\tilde{\Theta}_{m}(\bm{j},\bm{r}_{m})=\tanh\left[\frac{|\bm{j}-\bm{r}^{\mathrm{L}}_{m}|}{z^{\mathrm{L}}_{m}}\right],$ (21)
the size of vortices $z^{\mathrm{L}}_{m}$, and the core positions $\bm{r}^{\mathrm{L}}_{m}=(x^{\mathrm{L}}_{m},y^{\mathrm{L}}_{m})$ for $m=0,1$. Each phase vortex is given by
$\varphi^{\mathrm{L}}_{\mathrm{v}m}(\bm{j})=V^{\mathrm{L}}_{m}\arg[(j_{x}a-x^{\mathrm{L}}_{m})+i(j_{y}a-y^{\mathrm{L}}_{m})],$ (22)
where $V^{\mathrm{l}}$ is the number of windings for each phase vortex.
@@
Figure 5: Evaluation of the first and second harmonics Josephson current with two vortices. (a,b) $\bar{I}_{1}$ and (c,d) $\bar{J}_{1}$ for the $\bm{r}^{\mathrm{L}}_{1}=(x^{\mathrm{L}}_{1},y^{\mathrm{L}}_{1})$ space at (a,c) $(V^{\mathrm{L}}_{0},V^{\mathrm{L}}_{1})=(+1,+1)$ and (b,d) $(V^{\mathrm{L}}_{0},V^{\mathrm{L}}_{1})=(+1,-1)$. The color bars indicate the amplitude of (a,c) $\bar{I}_{1}(\bm{r}^{\mathrm{L}}_{0})$ and (b,d) $\bar{J}_{1}(\bm{r}^{\mathrm{L}}_{0})$. We set the core positions as $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})=(-30a,0.5a)$ and the vortex size as $z^{\mathrm{L}}_{1}=10a$ ($N^{\mathrm{L}}_{x}=N^{\mathrm{R}}_{x}=45$ and $N_{y}=30$). The Gray line indicates the scanning site in Fig. 4.